OpenStudy (anonymous):

solve by any method...s h o w w o r k 3y=4x+7 -4x-4y=28

OpenStudy (anonymous):

3y=4x+7............(1) -4x-4y=28 -4y = 4x + 28...........(2) (1) - (2):- 7y = -21 y = -3 plug y = -3 into (1):- -9 = 4x + 7 4x = -16 x = -4 solution x=-4, y=-3